OpenStudy (anonymous):

Evaluate cos(2Arctan(3/4) ) I know the answer is 7/25 but i'm not sure how to get the answer

OpenStudy (phi):

this can be confusing but we we start with tan A = 3/4 and take the arc tangent of both sides (this "undoes" the tangent), we get A = arctan(3/4) so we see that A = arctan(3/4) corresponds to tan A = 3/4 next, we see we want to find cos (2 * arctan(3/4) ) using A = arctan(3/4), this is asking for cos (2A)

OpenStudy (phi):

How do we find the cosine of 2A ? This will take multiple steps We start with A = arctan(3/4) and as noted above, this means tan A = 3/4 draw a triangle: